OpenStudy (anonymous):

Help!!!!!!!!!! URGENT!!!!!!!!!!! Jasmine estimated that her call would take 15 minutes but it actually took 12 minutes. What was her percent of error?

OpenStudy (anonymous):

i will give medal

OpenStudy (anonymous):

http://www.ehow.com/how_5127571_percent-error.html

jimthompson5910 (jim_thompson5910):

Percent Error = (|Estimate - Actual|/Actual) * 100 Percent Error = (|15-12|/12) * 100 Percent Error = (|3|/12) * 100 Percent Error = (3/12) * 100 Percent Error = (1/4) * 100 Percent Error = 100/4 Percent Error = 25%

OpenStudy (anonymous):

THANK U SOOOOOOOOOOO MUCH!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

you're welcome